OpenStudy (anonymous):

Can somebody help me please? (: Simplify: 1. 8/2s^-3 2. 6a^-1 c^-3/d^0 3. 12^0 t^7 u^-11

OpenStudy (anonymous):

https://www.mathway.com/ <--- That helps me. :)

OpenStudy (anonymous):

im possibly wrong but i think the first is -64

OpenStudy (anonymous):

\[\frac{ 8 }{ 2s^{-3} } = \frac{ 4 }{ s^{-3} } = 4s^3\]

OpenStudy (anonymous):

Just go by this rule: \[\frac{ 1 }{ a^{-b} } = a^b\]